Given f(x) = log5 x , find f(1/25)
here are the choices:
a. 5
b. -5
c. 1/5
d. -2

Answer by nerdybill(7384)

.
f(x) = log5 x
f(1/25) = log5 (1/25)
f(1/25) = log5(1) - log5(25)
f(1/25) = log5(5^0) - log5(5^2)
f(1/25) = 0 - 2
f(1/25) = -2
.
Answer: d.